On January 1, 2021, Weaver Corporation purchased a patent for $225,000. The remaining legal life is 20 years, but the company estimates the patent will be useful for only six more years. In January 2023, the company incurred legal fees of $45,000 in successfully defending a patent infringement suit. The successful defense did not change the company’s estimate of useful life. Weaver Corporation’s year-end is December 31.

What is the balance in the Patent account at the end of 2023?
Answer #1

It is assumed accumulated amortization is not used.

Amortization per year = $225000/6 = $37500

Value of Patent as on Jan 1, 2023 = $225000 - 37500x2 = $150000

Patent Value after defending patent = $150000 + 45000 = $195000
Amortization = $195000/4 = $48750

Balance in Patent at end of 2023 = $195000 - 48750 = $146250